Koontz Company manufactures a number of products. The standards relating to one of these products are shown below, along with actual cost data for May.

    Standard Cost per Unit         Actual Cost per Unit
Direct materials:                          
Standard: 1.80 feet at $1.20 per foot   $  
2.16
                  
Actual: 1.75 feet at $1.60 per foot                     $   2.80
Direct labor:                          
Standard: 0.90 hours at $16.00 per hour     
14.40
                  
Actual: 0.95 hours at $15.40 per hour                        14.63
Variable overhead:                          
Standard: 0.90 hours at $5.60 per hour      5.04                    
Actual: 0.95 hours at $5.00 per hour                        4.75
Total cost per unit   $  
21.60
             $   22.18
Excess of actual cost over standard cost per unit            $   0.58           
The production superintendent was pleased when he saw this report and commented: “This $0.58 excess cost is well within the 5 percent limit management has set for acceptable variances. It's obvious that there's not much to worry about with this product."

Actual production for the month was 10,500 units. Variable overhead cost is assigned to products on the basis of direct labor-hours. There were no beginning or ending inventories of materials.
Required:
1. Compute the following variances for May:
a. Materials price and quantity variances.
b. Labor rate and efficiency variances.
c. Variable overhead rate and efficiency variances.

2. How much of the $0.58 excess unit cost is traceable to each of the variances computed in (1) above.
3. How much of the $0.58 excess unit cost is traceable to apparent inefficient use of labor time?
Answer #1

Material Price Variance = (Standard Price – Actual Price)*Actual Quantity

= (1.20-1.60)*10500*1.8

= $7,560 U

Material Quantity Variance = (Standard Quantity – Actual Quantity)*Standard Price

= (1.80-1.75)*10,500*1.20

= $630 F

b.Labor Rate variance = (Standard Rate – Actual Rate)*Actual Hours

= (16-15.40)*0.95*10,500

= $5,985 F

Efficiency = (Standard Hours – Actual Hours)*Standard Rate

= (0.9-0.95)*10,500*16

= $8,400 U

c.Variable overhead Rate Variance = (Standard Rate – Actual Rate)*Actual Hours

=(5.60-5)*10,500*0.95

= $5,985 F

Efficiency = (0.9-0.95)*10,500*5.60

= $2,940 U

2.Direct Material Price = 7560/10500 = $0.72 U

Quantity = $0.06 F

Total = 0.66 U

Labor Rate = 0.57 F

Efficiency = 0.8 U

Total = 0.23 U

Variable overhead Rate = 0.57 F

Efficiency= 0.28 U

Total = 0.29 F

3.

portion due to inefficiency

Labor efficiency variance = 0.8 U

Variable overhead efficiency variance = 0.28 U

Total = 1.08 U